# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 5-shot | 12.22 | 84.62 | 21.36 | 90.91 | 20.41 | 33.33 | 18.23 | 99.11 | ||||||||||
| Gemma2 2B IT | 0-shot | 25.56 | 95.83 | 40.35 | 95.45 | 23.86 | 38.18 | 39.27 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 55.56 | 92.59 | 69.44 | 81.82 | 31.03 | 45.0 | 57.22 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 62.22 | 93.33 | 74.67 | 81.82 | 34.62 | 48.65 | 61.66 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 25.56 | 88.46 | 39.66 | 86.36 | 22.09 | 35.19 | 37.42 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 41.11 | 90.24 | 56.49 | 81.82 | 25.35 | 38.71 | 47.6 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 18.89 | 85.0 | 30.91 | 86.36 | 20.65 | 33.33 | 32.12 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 8.89 | 100.0 | 16.33 | 100.0 | 21.15 | 34.92 | 25.62 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 13.33 | 70.59 | 22.43 | 77.27 | 17.89 | 29.06 | 25.74 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 100.0 | 19.64 | 32.84 | 16.42 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 82.22 | 87.06 | 84.57 | 50.0 | 40.74 | 44.9 | 64.73 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 5.56 | 71.43 | 10.31 | 90.91 | 19.05 | 31.5 | 20.9 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 67.78 | 88.41 | 76.73 | 63.64 | 32.56 | 43.08 | 59.9 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 5.56 | 83.33 | 10.42 | 90.91 | 19.8 | 32.52 | 14.31 | 95.54 | |||||||||||
| DSM | Gemma2 2B IT | 5-shot | 25.93 | 70.0 | 37.84 | 77.78 | 41.18 | 53.85 | 30.56 | 97.78 | ||||||||||
| Gemma2 2B IT | 0-shot | 40.74 | 64.71 | 50.0 | 66.67 | 42.86 | 52.17 | 51.09 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 44.44 | 66.67 | 53.33 | 66.67 | 44.44 | 53.33 | 53.33 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 77.78 | 61.76 | 68.85 | 27.78 | 45.45 | 34.48 | 51.67 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 44.44 | 70.59 | 54.55 | 72.22 | 46.43 | 56.52 | 55.53 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 70.37 | 65.52 | 67.86 | 44.44 | 50.0 | 47.06 | 57.46 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 3.7 | 100.0 | 7.14 | 100.0 | 40.91 | 58.06 | 32.6 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 25.93 | 58.33 | 35.9 | 72.22 | 39.39 | 50.98 | 43.44 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 81.48 | 70.97 | 75.86 | 50.0 | 64.29 | 56.25 | 66.06 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 83.33 | 35.71 | 50.0 | 25.0 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 81.48 | 68.75 | 74.58 | 44.44 | 61.54 | 51.61 | 63.09 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 48.15 | 59.09 | 53.06 | 50.0 | 39.13 | 43.9 | 48.48 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 37.04 | 71.43 | 48.78 | 77.78 | 45.16 | 57.14 | 52.96 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 14.81 | 57.14 | 23.53 | 72.22 | 41.94 | 53.06 | 25.53 | 84.44 | |||||||||||
| AL | Gemma2 2B IT | 5-shot | 18.18 | 57.14 | 27.59 | 84.21 | 47.06 | 60.38 | 43.98 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 31.82 | 63.64 | 42.42 | 78.95 | 50.0 | 61.22 | 51.82 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 27.27 | 75.0 | 40.0 | 89.47 | 51.52 | 65.38 | 52.69 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 63.64 | 66.67 | 65.12 | 63.16 | 60.0 | 61.54 | 63.33 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 95.45 | 60.0 | 73.68 | 26.32 | 83.33 | 40.0 | 56.84 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 4.55 | 100.0 | 8.7 | 100.0 | 47.5 | 64.41 | 36.55 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 9.09 | 50.0 | 15.38 | 89.47 | 45.95 | 60.71 | 38.05 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 45.45 | 50.0 | 47.62 | 47.37 | 42.86 | 45.0 | 46.31 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 81.82 | 60.0 | 69.23 | 36.84 | 63.64 | 46.67 | 57.95 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 4.55 | 33.33 | 8.0 | 89.47 | 44.74 | 59.65 | 33.82 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 27.27 | 66.67 | 38.71 | 84.21 | 50.0 | 62.75 | 50.73 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 0.0 | 0.0 | 0.0 | 94.74 | 45.0 | 61.02 | 30.51 | 100.0 | |||||||||||
| AD | Gemma2 2B IT | 5-shot | 0.0 | 0.0 | 0.0 | 100.0 | 47.37 | 64.29 | 32.14 | 100.0 | ||||||||||
| Gemma2 2B IT | 0-shot | 45.0 | 64.29 | 52.94 | 72.22 | 54.17 | 61.9 | 57.42 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 70.0 | 60.87 | 65.12 | 50.0 | 60.0 | 54.55 | 59.83 | 100.0 | |||||||||||
| Qwen2 7B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 80.0 | 69.57 | 74.42 | 61.11 | 73.33 | 66.67 | 70.54 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 30.0 | 54.55 | 38.71 | 72.22 | 48.15 | 57.78 | 48.24 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 10.0 | 66.67 | 17.39 | 94.44 | 48.57 | 64.15 | 40.77 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 25.0 | 83.33 | 38.46 | 94.44 | 53.12 | 68.0 | 53.23 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 5.0 | 100.0 | 9.52 | 100.0 | 48.65 | 65.45 | 37.49 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 20.0 | 66.67 | 30.77 | 88.89 | 50.0 | 64.0 | 47.38 | 100.0 | |||||||||||
| AC | Gemma2 2B IT | 5-shot | 22.22 | 66.67 | 33.33 | 86.84 | 49.25 | 62.86 | 32.06 | 98.8 | ||||||||||
| Gemma2 2B IT | 0-shot | 15.56 | 87.5 | 26.42 | 97.37 | 49.33 | 65.49 | 45.95 | 100.0 | |||||||||||
| Gemma2 9B IT | 5-shot | 15.56 | 100.0 | 26.92 | 100.0 | 50.0 | 66.67 | 46.79 | 100.0 | |||||||||||
| Gemma2 9B IT | 0-shot | 26.67 | 54.55 | 35.82 | 73.68 | 47.46 | 57.73 | 31.18 | 97.59 | |||||||||||
| Qwen2 7B IT | 5-shot | 15.56 | 77.78 | 25.93 | 94.74 | 48.65 | 64.29 | 45.11 | 100.0 | |||||||||||
| Qwen2 7B IT | 0-shot | 48.89 | 56.41 | 52.38 | 55.26 | 47.73 | 51.22 | 51.8 | 100.0 | |||||||||||
| SaigaLlama3 8B | 5-shot | 13.33 | 100.0 | 23.53 | 100.0 | 49.35 | 66.09 | 44.81 | 100.0 | |||||||||||
| SaigaLlama3 8B | 0-shot | 46.67 | 67.74 | 55.26 | 73.68 | 53.85 | 62.22 | 58.74 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 37.78 | 65.38 | 47.89 | 76.32 | 50.88 | 61.05 | 54.47 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 17.78 | 53.33 | 26.67 | 81.58 | 45.59 | 58.49 | 42.58 | 100.0 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 51.11 | 57.5 | 54.12 | 55.26 | 48.84 | 51.85 | 52.98 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 88.89 | 54.79 | 67.8 | 13.16 | 50.0 | 20.83 | 44.31 | 100.0 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 22.22 | 62.5 | 32.79 | 73.68 | 50.0 | 59.57 | 30.79 | 86.75 |

### v3

@@ Appendix B Full Results for Models
| DE | Gemma2 2B IT | 5-shot | 82.38 1.12 | 10.44 0.89 | 18.53 1.41 | 20.08 0.16 | 90.91 0.00 | 32.90 0.22 | 17.14 0.54 | 99.11 0.00 | ||||||||||
| Gemma2 2B IT | 0-shot | 95.83 0.00 | 25.56 0.00 | 40.35 0.00 | 23.86 0.00 | 95.45 0.00 | 38.18 0.00 | 39.27 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 92.80 0.11 | 57.33 0.89 | 70.88 0.72 | 31.92 0.44 | 81.82 0.00 | 45.92 0.46 | 58.40 0.59 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 93.33 0.00 | 62.22 0.00 | 74.67 0.00 | 34.62 0.00 | 81.82 0.00 | 48.65 0.00 | 61.66 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 88.80 0.17 | 26.44 0.44 | 40.75 0.55 | 22.30 0.10 | 86.36 0.00 | 35.45 0.13 | 38.10 0.34 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 92.33 1.05 | 42.89 0.89 | 58.57 1.04 | 26.78 0.72 | 85.45 1.82 | 40.79 1.04 | 49.68 1.04 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 85.00 0.00 | 18.89 0.00 | 30.91 0.00 | 20.65 0.00 | 86.36 0.00 | 33.33 0.00 | 32.12 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 100.00 0.00 | 8.00 0.44 | 14.81 0.76 | 20.99 0.08 | 100.00 0.00 | 34.70 0.11 | 24.76 0.43 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 73.07 1.24 | 15.11 0.89 | 25.04 1.30 | 18.20 0.15 | 77.27 0.00 | 29.46 0.20 | 27.25 0.75 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 19.64 0.00 | 100.00 0.00 | 32.84 0.00 | 16.42 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 87.06 0.00 | 82.22 0.00 | 84.57 0.00 | 40.74 0.00 | 50.00 0.00 | 44.90 0.00 | 64.73 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 71.43 0.00 | 5.56 0.00 | 10.31 0.00 | 19.05 0.00 | 90.91 0.00 | 31.50 0.00 | 20.90 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 88.54 0.07 | 68.67 0.44 | 77.35 0.31 | 33.18 0.31 | 63.64 0.00 | 43.62 0.27 | 60.48 0.29 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 85.24 0.95 | 6.44 0.44 | 11.98 0.78 | 20.59 0.40 | 94.55 1.82 | 33.82 0.65 | 15.27 0.48 | 96.25 0.36 | |||||||||||
| DSM | Gemma2 2B IT | 5-shot | 54.00 8.00 | 11.11 7.41 | 17.89 9.97 | 39.45 0.86 | 86.67 4.44 | 54.16 0.16 | 32.97 1.20 | 99.56 0.89 | ||||||||||
| Gemma2 2B IT | 0-shot | 64.71 0.00 | 40.74 0.00 | 50.00 0.00 | 42.86 0.00 | 66.67 0.00 | 52.17 0.00 | 51.09 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 64.24 1.21 | 29.63 7.41 | 40.14 6.60 | 41.83 1.31 | 75.56 4.44 | 53.74 0.21 | 46.94 3.20 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 62.64 0.44 | 80.74 1.48 | 70.54 0.85 | 49.09 1.82 | 27.78 0.00 | 35.47 0.49 | 53.01 0.67 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 70.59 0.00 | 44.44 0.00 | 54.55 0.00 | 46.43 0.00 | 72.22 0.00 | 56.52 0.00 | 55.53 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 65.52 0.00 | 70.37 0.00 | 67.86 0.00 | 50.00 0.00 | 44.44 0.00 | 47.06 0.00 | 57.46 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 3.70 0.00 | 7.14 0.00 | 40.91 0.00 | 100.00 0.00 | 58.06 0.00 | 32.60 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 55.30 1.52 | 22.96 1.48 | 32.44 1.73 | 38.47 0.46 | 72.22 0.00 | 50.20 0.39 | 41.32 1.06 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 69.19 0.89 | 81.48 0.00 | 74.83 0.51 | 62.09 1.10 | 45.56 2.22 | 52.54 1.85 | 63.69 1.18 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 35.71 0.00 | 83.33 0.00 | 50.00 0.00 | 25.00 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 68.75 0.00 | 81.48 0.00 | 74.58 0.00 | 61.54 0.00 | 44.44 0.00 | 51.61 0.00 | 63.09 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 59.09 0.00 | 48.15 0.00 | 53.06 0.00 | 39.13 0.00 | 50.00 0.00 | 43.90 0.00 | 48.48 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 61.34 5.04 | 37.04 0.00 | 46.12 1.33 | 40.46 2.35 | 64.44 6.67 | 49.69 3.73 | 47.90 2.53 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 57.14 0.00 | 14.81 0.00 | 23.53 0.00 | 41.94 0.00 | 72.22 0.00 | 53.06 0.00 | 25.53 0.00 | 84.44 0.00 | |||||||||||
| AL | Gemma2 2B IT | 5-shot | 57.14 0.00 | 18.18 0.00 | 27.59 0.00 | 47.06 0.00 | 84.21 0.00 | 60.38 0.00 | 43.98 0.00 | 100.00 0.00 | ||||||||||
| Gemma2 2B IT | 0-shot | 63.64 0.00 | 31.82 0.00 | 42.42 0.00 | 50.00 0.00 | 78.95 0.00 | 61.22 0.00 | 51.82 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 75.00 0.00 | 27.27 0.00 | 40.00 0.00 | 51.52 0.00 | 89.47 0.00 | 65.38 0.00 | 52.69 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 66.67 0.00 | 63.64 0.00 | 65.12 0.00 | 60.00 0.00 | 63.16 0.00 | 61.54 0.00 | 63.33 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 61.33 2.67 | 27.27 0.00 | 37.74 0.48 | 48.71 0.65 | 80.00 2.11 | 60.55 1.10 | 49.15 0.79 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 60.00 0.00 | 95.45 0.00 | 73.68 0.00 | 83.33 0.00 | 26.32 0.00 | 40.00 0.00 | 56.84 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 4.55 0.00 | 8.70 0.00 | 47.50 0.00 | 100.00 0.00 | 64.41 0.00 | 36.55 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 36.67 6.67 | 5.45 1.82 | 9.48 2.95 | 44.98 0.48 | 89.47 0.00 | 59.86 0.43 | 34.67 1.69 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 50.00 0.00 | 45.45 0.00 | 47.62 0.00 | 42.86 0.00 | 47.37 0.00 | 45.00 0.00 | 46.31 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 45.00 0.00 | 94.74 0.00 | 61.02 0.00 | 30.51 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 60.00 0.00 | 81.82 0.00 | 69.23 0.00 | 63.64 0.00 | 36.84 0.00 | 46.67 0.00 | 57.95 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 33.33 0.00 | 4.55 0.00 | 8.00 0.00 | 44.74 0.00 | 89.47 0.00 | 59.65 0.00 | 33.82 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 66.67 0.00 | 27.27 0.00 | 38.71 0.00 | 50.00 0.00 | 84.21 0.00 | 62.75 0.00 | 50.73 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 43.87 0.56 | 90.53 2.11 | 59.10 0.96 | 29.55 0.48 | 100.00 0.00 | |||||||||||
| AD | Gemma2 2B IT | 5-shot | 0.00 0.00 | 0.00 0.00 | 0.00 0.00 | 47.37 0.00 | 100.00 0.00 | 64.29 0.00 | 32.14 0.00 | 100.00 0.00 | ||||||||||
| Gemma2 2B IT | 0-shot | 64.29 0.00 | 45.00 0.00 | 52.94 0.00 | 54.17 0.00 | 72.22 0.00 | 61.90 0.00 | 57.42 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 60.87 0.00 | 70.00 0.00 | 65.12 0.00 | 60.00 0.00 | 50.00 0.00 | 54.55 0.00 | 59.83 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 67.25 1.16 | 80.00 0.00 | 73.07 0.68 | 71.81 0.76 | 56.67 2.22 | 63.33 1.67 | 68.20 1.17 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 54.55 0.00 | 30.00 0.00 | 38.71 0.00 | 48.15 0.00 | 72.22 0.00 | 57.78 0.00 | 48.24 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 50.91 1.82 | 30.00 0.00 | 37.74 0.48 | 46.55 0.80 | 67.78 2.22 | 55.19 1.29 | 46.47 0.89 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 13.33 26.67 | 2.00 4.00 | 3.48 6.96 | 46.47 1.05 | 94.44 0.00 | 62.28 0.93 | 32.88 3.94 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 66.67 0.00 | 10.00 0.00 | 17.39 0.00 | 48.57 0.00 | 94.44 0.00 | 64.15 0.00 | 40.77 0.00 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 83.33 0.00 | 25.00 0.00 | 38.46 0.00 | 53.12 0.00 | 94.44 0.00 | 68.00 0.00 | 53.23 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 100.00 0.00 | 5.00 0.00 | 9.52 0.00 | 48.65 0.00 | 100.00 0.00 | 65.45 0.00 | 37.49 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 66.67 0.00 | 20.00 0.00 | 30.77 0.00 | 50.00 0.00 | 88.89 0.00 | 64.00 0.00 | 47.38 0.00 | 100.00 0.00 | |||||||||||
| AC | Gemma2 2B IT | 5-shot | 72.00 2.67 | 24.00 0.89 | 36.00 1.33 | 50.45 0.60 | 88.95 1.05 | 64.38 0.76 | 33.46 0.70 | 98.80 0.00 | ||||||||||
| Gemma2 2B IT | 0-shot | 97.50 5.00 | 15.56 0.00 | 26.82 0.20 | 49.87 0.27 | 99.47 1.05 | 66.43 0.47 | 46.63 0.34 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 5-shot | 68.00 16.00 | 13.78 0.89 | 22.84 2.04 | 47.26 1.37 | 91.58 4.21 | 62.34 2.16 | 42.59 2.10 | 100.00 0.00 | |||||||||||
| Gemma2 9B IT | 0-shot | 54.55 0.00 | 26.67 0.00 | 35.82 0.00 | 47.46 0.00 | 73.68 0.00 | 57.73 0.00 | 31.18 0.00 | 97.59 0.00 | |||||||||||
| Qwen2 7B IT | 5-shot | 84.89 3.56 | 26.22 5.33 | 39.85 6.96 | 52.08 1.72 | 94.74 0.00 | 67.20 1.46 | 53.52 4.21 | 100.00 0.00 | |||||||||||
| Qwen2 7B IT | 0-shot | 55.93 0.24 | 52.44 1.78 | 54.11 0.87 | 47.55 0.09 | 51.05 2.11 | 49.22 1.00 | 51.67 0.07 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 5-shot | 100.00 0.00 | 13.33 0.00 | 23.53 0.00 | 49.35 0.00 | 100.00 0.00 | 66.09 0.00 | 44.81 0.00 | 100.00 0.00 | |||||||||||
| SaigaLlama3 8B | 0-shot | 66.05 0.85 | 46.67 0.00 | 54.69 0.29 | 53.12 0.36 | 71.58 1.05 | 60.98 0.62 | 57.84 0.45 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 5-shot | 54.20 0.30 | 88.89 0.00 | 67.34 0.23 | 45.56 2.22 | 11.05 1.05 | 17.78 1.52 | 42.56 0.88 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 0.4 | 0-shot | 67.48 1.05 | 37.78 0.00 | 48.43 0.27 | 51.55 0.34 | 78.42 1.05 | 62.21 0.58 | 55.32 0.43 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 5-shot | 90.67 18.67 | 8.89 4.44 | 15.33 5.67 | 47.12 0.76 | 96.32 7.37 | 63.22 2.37 | 39.28 1.65 | 100.00 0.00 | |||||||||||
| Vikhr 7B IT 5.4 | 0-shot | 57.50 0.00 | 51.11 0.00 | 54.12 0.00 | 48.84 0.00 | 55.26 0.00 | 51.85 0.00 | 52.98 0.00 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 5-shot | 52.75 1.02 | 80.00 4.44 | 63.56 2.12 | 40.00 5.00 | 15.26 1.05 | 21.94 0.56 | 42.75 0.78 | 100.00 0.00 | |||||||||||
| VikhrGemma 2B IT | 0-shot | 64.26 0.88 | 24.00 0.89 | 34.94 1.08 | 50.00 0.00 | 73.68 0.00 | 59.57 0.00 | 31.51 0.36 | 87.71 0.48 |